OpenStudy (anonymous):

e^t is a solution of ty'' - (2t+1)y' + (t+1)y = 0 what is the general solution

OpenStudy (anonymous):

\[y_1=e^t\] let \(y_2=vy_1\) and find \(v\)

OpenStudy (anonymous):

are u familiar with Reduction of Order method?

OpenStudy (anonymous):

If you know one solution \(y_1\), you can find the other one by substituting \(y_2 = vy_1\) in the homogenous equation

OpenStudy (anonymous):

my v term didnt drop

OpenStudy (anonymous):

so tv''e^t + v'e^t + ve^t = 0

OpenStudy (experimentx):

that should drop ...  equate this term to zero ... your v drops.